Zeldris:

A satellite orbiting Earth at an orbital radius r has a velocity v. Which represents the velocity if the satellite is moved to an orbital radius of 4r?

Alexis1415:

Well let's see, V=GM/4r, V=1/4x GM/r V=1/4 x V The orbital radius is 4r it would be V=v over \[\sqrt{4}\]

Zeldris:

thanks

Alexis1415:

@zeldris wrote:
thanks
np c: